# Self Pre-training with Topology- and Spatiality-aware Masked Autoencoders for 3D Medical Image Segmentation

### v2

@@ Abstract
P. Gu is with the Department of Computer Science, University of Texas Rio Grande Valley, Edinburg, TX 78539, USA. H. Li is with Department of Mathematical Sciences, The University of Texas at Dallas, Richardson, TX 75080, USA. Y. Zhang, C. Wang, and D. Chen are with the Department of Computer Science and Engineering, University of Notre Dame, Notre Dame, IN 46556, USA. E-mail: pengfei.gu01@utrgv.edu, dchen@nd.edu
Index Terms—  Self-supervised Learning, Masked Autoencoders, Topology, Spatiality, 3D Medical Image Segmentation
1 Introduction
@@
Accurate segmentation of medical images is critical for medical analysis and applications such as diagnosis, treatment planning, and research. While many deep learning (DL) models, such as nnU-Net [1], nnFormer [2], and UNETR++ [3], have demonstrated impressive performances in 3D medical image segmentation, such methods still face several key challenges. A primary challenge is the scarcity of high-quality labeled medical images for model training, due to high costs and expertise needed for data collection and annotation. Another challenge is annotation errors, as labeling 3D medical images can be very time-consuming and error-prone.
Self-supervised learning (SSL), a technique that leverages pre-text tasks to derive useful visual representations from unlabeled data, offers a promising avenue to combat the challenge of label scarcity. One representative methodology for SSL is Masked Autoencoders (MAEs) [4]. Specifically, MAE learns to reconstruct the missing pixels after randomly masking a certain fraction (e.g., 75%) of patches of the input images. This approach has been successfully deployed in various applications, including in image domain [5, 6], video domain [7, 8, 9], and language domain [10, 11].
@@
In the medical image segmentation area, MAE pre-training has also been found to be effective (e.g., UNETR + MAE [12]). Although simple and effective, there are still several limitations. First, geometric shape information (i.e., contextual information on the overall shapes of objects), which is critical for improving segmentation performance, is not captured well (e.g., see Fig. 1). Second, global spatial information is not well explored since the focus has been on reconstructing information from the masked local sub-volumes, possibly neglecting the global context information of the target objects as a whole. Third, the MAE pre-training strategy (i.e., learning representations by reconstructing missing patches from masked image input) is not exploited well with various common medical image segmentation architectures, e.g., those based on convolutional neural networks (CNNs) or hybrid models. This is primarily because MAE was developed using the Vision Transformer (ViT) [13] architecture, potentially restricting its adaptability and effectiveness with other architectures.
@@
To address these limitations, we propose a novel extension of MAEs for self pre-training for 3D medical image segmentation. (I) We extract geometric shape information by exploiting multi-scale topological features (e.g., connected components, cycles/loops, and voids). Our method utilizes cubical complexes [14, 15] to compute topological signatures of both the input and reconstructed volumes, and employs an optimal transport distance (the 2-Wasserstein distance [16]) to derive a new topological loss. Our topology-aware loss is fully differentiable, computationally efficient, can be added to any neural network, and is applicable to 2D/3D images. (II) We propose a pre-text task to predict the positions of multiple key points of crops, enabling the model to aggregate spatial information. Specifically, our method predicts the positions of nine points (the center and eight corners) of a 3D crop in the input volume. By learning where the crops are located in the input volume, the model can capture global spatial information. (III) We extend the MAE pre-training strategy to a hybrid state-of-the-art (SOTA) medical image segmentation architecture, UNETR++ [3], and co-pretrain UNETR++ alongside the ViT. Specifically, masked crops are processed separately by both ViT and UNETR++ to reconstruct the associated missing patches. Reconstruction consistency loss and spatial consistency loss (derived from the pre-text task) are employed to connect the two different types of architectures in pre-training, enhancing their representation learning capability.
@@
Our main contributions are summarized as follows:
  1. 1.
    We propose a new topological loss and introduce a pre-text task for MAEs to learn geometric shape and spatial information.
  2. 2.
    We extend the MAE pre-training strategy to a hybrid SOTA medical image segmentation architecture and co-pretrain it alongside ViT.
  3. 3.
    We develop a fine-tuned model for downstream segmentation tasks, and demonstrate the effectiveness of our new approach on five public 3D segmentation datasets.
@@
Fig. 1: Illustrating the effect of our proposed topological loss. Left column: raw image examples; middle column: reconstructed images with the mean squared error (MSE) loss [12]; right column: reconstructed images with a combination of the MSE and proposed topological losses. (a) Image examples of the Synapse multi-organ CT segmentation dataset [17]. (b) Image examples of spleen segmentation on the Medical Segmentation Decathlon (MSD) dataset [18]. The proposed topological loss helps to effectively recover objects with correct topology.
2 Related Work
In this section, we review recent methods pertinent to our work, including SSL, MAE, topology-based DL methods, and related DL-based medical image segmentation architectures.
Self-supervised Learning (SSL). SSL is widely recognized as an effective way to alleviate the annotation burden as it can learn useful visual representations from unlabeled data. Most known SSL methods can be categorized as either generative or discriminative.
Generative SSL methods commonly learn image representations by generating images pixel-for-pixel, using training images as supervision. The typical methods include image generation with GAN [19], super-resolution [20], image inpainting [21], image decoupling [22], and image colorization [23]. But, pixel-for-pixel generation incurs high computation costs and is difficult to learn more abstract latent representations.
Discriminative SSL methods often learn image representations by constructing discriminative pre-text tasks. Common pre-text tasks include image jigsaw puzzle [24], geometric transformations [25], and clustering [26]. Recently, contrastive learning, a new discriminative SSL method, has shown strong performance. SimCLR [27] and MoCo [28] are two typical contrastive learning methods that seek to learn representations by attracting positive pairs and repulsing negative pairs. However, these methods rely on a large number of negative samples in order to work well. BYOL [29] improved contrastive learning by directly predicting the representation output from another view and discarding negative pairs, yielding SOTA performance. SimSiam [30] further showed that stop-gradient plays an essential role in the learning stability of siamese neural networks. Since positive pairs in BYOL come from the same image, feature diversity from different images is ignored.
While our approach is SSL, it targets self pre-training paradigms for 3D medical image segmentation.
Masked Autoencoders (MAEs). MAEs [4] are a type of denoising autoencoders [31] that learn representations by reconstructing the original input from masked input. The introduction of masked language modeling in BERT [32] gave a transformative impact on natural language processing (NLP). Recently, several studies applied MAEs to medical image analysis [33, 6, 12, 34, 35, 36]. In [12], it showed that directly applying MAEs to medical datasets yields superior performance over its counterpart pre-trained on ImageNet [37]. In [6], it exhibited that MAEs can help 3D medical image analysis, besides natural images. In [34, 35], it proved that adding an additional global loss on top of a local loss can make representations more semantically meaningful. In [33], it directly applied MAEs to medical image classification. A self-distillation loss was used to improve MAE pre-training [36].
Different from these known methods, our approach emphasizes extending the MAE pre-training strategy to a hybrid medical image segmentation architecture, enhancing its ability to learn geometric shape and global spatial information.
Topology-based Deep Learning Methods. Topology is often referred to as encoding the overall shapes of data. Numerous works aim to leverage topological information in DL methods to tackle various biomedical image analysis tasks.
For segmentation tasks, pre-trained filters were leveraged to detect higher-order topological features of linear structures [38]. clDice [39] proposed an additional Dice loss for extracted skeleton structures, guaranteeing topology preservation. A persistent homology (PH) based loss was designed for DL architectures [40, 41], learning to segment with correct topology. DMT-loss [42] utilized discrete Morse theory to identify topologically critical structures for segmentation.
For classification tasks, a new topological regularizer was proposed [43] to train classifiers for classifying kidney and brain cell cancers. A unified model was presented that aggregates appearance, microenvironment, and topology for pathology image classification [44]. For reconstruction tasks, a multi-scale topological loss was proposed [45] to capture shape information for 3D reconstruction (i.e., reconstructing 3D objects from 2D images).
For other tasks, a method was introduced to detect topological handles for cardiac trabeculae restoration [46]. A topology-aware generative adversarial network was proposed to generate images with correct topology [47]. Topographical features were computed for clustering high dimensional categorical data [48]. A GO-term embedding based method was proposed for protein semantic similarity prediction [49].
Different from these known methods, our method computes a new topological loss on raw images. To our best knowledge, no known methods employ a topology-aware loss for self pre-training for 3D medical image segmentation.
DL Medical Image Segmentation Methods. Most DL medical image segmentation methods can be categorized as CNN-based, Transformer-based, or hybrid. Many CNN-based segmentation methods are extended from the standard U-Net architecture [50]; some typical such models include Attention U-Net [51], UNet++ [52], UNet 3+ [53], and nnU-Net [1].
Due to their ability to encode long-range dependencies, ViTs recently gained popularity. A main building block in the Transformer architecture is the self attention operation that models interactions among a sequence of image patches, thus learning global relationships. There are a few pure Transformer-based methods for medical image segmentation [54, 55].
The known hybrid methods combine convolutions and self-attention operations for better segmentation. Recent common hybrid methods include TransUNet [56], UNETR [57], MISSFormer [58], Swin UNETR [59], nnFormer [2], UNETR++ [3], TransBTS [60], CoTr [61], TransFuse [62], MedT [63], and ConvFormer [64].
In this work, our focus is not on proposing a new DL architecture for 3D medical image segmentation. Instead, we aim to construct a fine-tuned model on top of an existing SOTA architecture (i.e., UNETR++ [3]).
Fig. 2: An overview of our proposed pipeline. In the self pre-training stage, we randomly mask a fraction (e.g., 50%) of patches of the 3D crops. The masked crops are processed independently by both a ViT model and a UNETR++ model [3], which are pre-trained with our proposed topological loss, pre-text task that predicts the positions of 9 key points of the crops, and spatial and reconstruction consistency losses for learning geometric shape and global spatial information. In the fine-tuning stage, the pre-trained ViT encoder is complemented with the pre-trained UNETR++ model, which is then fine-tuned for the target segmentation task. For visual simplicity, this illustration uses images instead of sub-volumes (displayed in 2D), shows only 5 key points (the five red points in the original image in (a)) for the spatial loss, and omits the MSE reconstruction loss for the two branches in the self pre-training stage in (b).
3 Method
@@
In Subsection 3.1, we introduce the topological loss for extracting the topology of the input volume. In Subsection 3.2, we propose the pre-text task for capturing global spatial information. In Subsection 3.3, we present the spatial and reconstruction consistency losses for co-pretraining the ViT and UNETR++ models. In Subsection 3.4, we construct the fine-tuned architecture for the downstream segmentation tasks.
3.1 Capturing the Topology of Input Volumes
Given a 3D image $I$, we represent $I$ with a cubical complex $C$. Typically, the cubical complex $C$ takes each voxel of $I$ as an individual vertex and contains connectivity information on vertex neighbourhoods via edges, squares, and their higher-dimensional counterparts [14, 15]. In this work, we use persistent homology (PH) [65] to extract topological features of different dimensions from $C$, including connected components (00-D), cycles/loops ($1$-D), and voids ($2$-D). PH combines the homology of super-level sets by sweeping a threshold function through the entire real numbers. Specifically, for a threshold value $\tau\in\mathbb{R}$, a cubical complex is defined as:
$C^{(\tau)}:=\{x\in I\ |\ f(x)\geq\tau\},$ (1)
where $f(x)$ is the voxel value of $x$. When sweeping the threshold, the topology changes only at a finite number of values, $\tau_{1}\geq\tau_{2}\geq\cdots\geq\tau_{m-1}\geq\tau_{m}$, and we obtain a sequence of nested cubical complexes, $\emptyset\subseteq C^{(\tau_{1})}\subseteq C^{(\tau_{2})}\subseteq\cdots\subseteq C^{(\tau_{m-1})}\subseteq C^{(\tau_{m})}=I$, which forms the super-level set filtration. PH tracks topological features across all the complexes in this filtration, representing each feature as a tuple $(\tau_{i},\tau_{j})$ with $\tau_{i}\geq\tau_{j},$ indicating the cubical complex in which a feature appears and disappears, respectively. For example, a 00-D tuple $(\tau_{i},\tau_{j})$ represents a connected component that appears at threshold $\tau_{i}$ and disappears at threshold $\tau_{j}$. The tuples of the $k$-D ($0\leq k\leq 2$) features are saved in the $k$-th persistence diagram $D_{I}^{k}$, which is a multi-scale shape descriptor of all topological features of the 3D image $I$.
Comparing Persistence Diagrams. Given two persistence diagrams $D$ and $D^{\prime}$, we use the 2-Wasserstein distance [16] as a metric to measure their similarity or distance, defined as:
$W_{2}(D,D^{\prime}):=\left(\inf_{\eta:D\rightarrow D^{\prime}}\sum_{x\in D}||x-\eta(x)||_{\infty}^{2}\right)^{\frac{1}{2}},$ (2)
where $\eta(\cdot)$ denotes a bijection. Note that Eq. (2) can be solved by using an optimal transport algorithm [66], and we use cubical Ripser [15] to compute PHs from volumes.
Constructing Topological Loss. Given an input volume $I$ and a reconstructed volume $I^{\prime}$, our new topology-aware loss is defined as:
$\mathcal{L}_{topo}(I,I^{\prime})=\left(\sum_{i=0}^{2}(W_{2}(D_{I}^{i},D_{I^{\prime}}^{i}))^{2}\right)^{\frac{1}{2}},$ (3)
where $W_{2}(\cdot,\cdot)$ denotes the 2-Wasserstein distance, and $D_{I}^{i}$ and $D_{I^{\prime}}^{i}$ are the $i$-D persistence diagrams of $I$ and $I^{\prime}$, respectively. Note that our proposed topological loss differs from that in [45]. First, we extract topological features from 3D images, not from the segmentation. Second, we use the topological loss for self pre-training and medical image segmentation, not for 3D reconstruction tasks.
3.2 Exploiting Global Spatial Information
@@
Constructing Spatial Loss. We denote the 9 key points of a 3D crop as $(p_{1},p_{2}$, $\ldots,p_{9})$, where $p_{9}$ is for the crop center, and each $p_{i}=(x_{i},y_{i},z_{i})$. Given the ground truth (GT) and the prediction of the 9 positions, $P=(p_{1},p_{2},\ldots,p_{9})$ and $P^{\prime}=(p^{\prime}_{1},p^{\prime}_{2},\ldots,p^{\prime}_{9})$, the spatial loss is defined as:
$\mathcal{L}_{spa}(P,P^{\prime})=\mathcal{L}_{MSE}(P,P^{\prime}),$ (4)
where $\mathcal{L}_{MSE}$ is the MSE loss. Note that our spatial loss definition and implementation are different from those in [67].
3.3 Co-pretraining the ViT and UNETR++ Models
@@
Constructing Spatial Consistency Loss. Given predictions of 9 key point positions from the ViT and UNETR++ encoders, $P^{\prime}=(p^{\prime}_{1},p^{\prime}_{2},\ldots,p^{\prime}_{9})$ and $P^{\prime\prime}=(p^{\prime\prime}_{1},p^{\prime\prime}_{2},\ldots,p^{\prime\prime}_{9})$, the spatial consistency loss is defined as:
$\mathcal{L}_{spa-consis}(P^{\prime},P^{\prime\prime})=\mathcal{L}_{MSE}(P^{\prime},P^{\prime\prime}).$ (5)
The spatial consistency loss aligns the reconstructed images at the spatial level, enhancing the representation learning capability of both the ViT and UNETR++ models.
Constructing Reconstruction Consistency Loss. Given reconstructed volumes $I^{\prime}$ and $I^{\prime\prime}$ by the ViT and UNETR++ models, our reconstruction consistency loss function computes the MSE between the reconstructed volumes $I^{\prime}$ and $I^{\prime\prime}$, as:
$\mathcal{L}_{rec-consis}(I^{\prime},I^{\prime\prime})=\mathcal{L}_{MSE}(I^{\prime},I^{\prime\prime}).$ (6)
We compute this loss only on masked patches, similar to MAE in [4]. The reconstruction consistency loss aligns the reconstructed images at the image level, further enhancing the representation learning capability of both ViT and UNETR++.
@@
3.4 Constructing the Fine-tuned Architecture
@@
Table 1: Segmentation results of different methods on the Synapse CT dataset. The results of the baselines are from [3, 12]. “—” indicates that the results were not reported in the original papers. The reported values are the averages of 5 runs with different random seeds. The best results are marked in bold, and the second-best results are underlined. We calculate the $p$-values between the average performance of our method and the best performing baseline in both the metrics.
@@
[…]
$p$-values $<5e-2$ (Dice), $<1e-2$ (HD95)
4 Experiments and analysis
We conduct comprehensive experiments to validate the effectiveness of our proposed approach. In Subsection 4.1, we discuss the datasets and experimental setup. In Subsection 4.2, we describe the implementation details. In Subsection 4.3, we present the metrics for evaluating the performance of our method. In Subsection 4.4, we report the experimental results. In Subsection 4.5, we show the ablation study to examine the effects of different key components in our method.
4.1 Datasets and Experimental Setup

### v3

@@ Abstract
I Introduction
@@
Accurate segmentation of medical images is critical for medical analysis and applications such as diagnosis, treatment planning, and research. While many deep learning (DL) models (e.g.,  [1, 25]) have demonstrated impressive performances in medical image segmentation, such methods still face several key challenges. One challenge is the scarcity of high-quality labeled medical images for model training, due to high costs and expertise needed for data collection and annotation. Another challenge is annotation errors, as labeling 3D medical images can be very time-consuming and error-prone.
@@
Self-supervised learning (SSL), a technique that leverages pre-text tasks to derive useful visual representations from unlabeled data, offers a promising avenue to combat the challenge of label scarcity. One representative methodology for SSL is Masked Autoencoders (MAEs) [11]. Specifically, MAE learns to reconstruct the missing pixels after randomly masking a certain fraction (e.g., 75%) of patches of the input images. In the medical image segmentation area, MAE pre-training has also been found to be effective (e.g., UNETR + MAE [28]). Although simple and effective, there are still several limitations. First, geometric shape information (i.e., contextual information on the overall shapes of objects), which is critical for improving segmentation performance, is not captured well (e.g., see Fig. 1). Second, global spatial information is not well explored since the focus has been on reconstructing information from the masked local sub-volumes, possibly neglecting the global context information of the target objects as a whole. Third, the MAE pre-training strategy (i.e., learning representations by reconstructing missing patches from masked image input) is not exploited well with various common medical image segmentation architectures, e.g., those based on convolutional neural networks (CNNs) or hybrid models. This is primarily because MAE was developed using the Vision Transformer (ViT) [6] architecture, potentially restricting its adaptability and effectiveness with other architectures.
@@
To address these limitations, we propose a novel extension of MAEs for self pre-training for 3D medical image segmentation. (I) We extract geometric shape information by exploiting multi-scale topological features (e.g., connected components, cycles/loops, and voids). Our method utilizes cubical complexes [14] to compute topological signatures of both the input and reconstructed volumes, and employs an optimal transport distance (the 2-Wasserstein distance) to derive a new topological loss. Our topology-aware loss is fully differentiable, computationally efficient, can be added to any neural network, and is applicable to 2D/3D images. (II) We propose a pre-text task to predict the positions of multiple key points of crops, enabling the model to aggregate spatial information. Specifically, our method predicts the positions of nine points (the center and eight corners) of a 3D crop in the input volume. By learning where the crops are located in the input volume, the model can capture global spatial information. (III) We extend the MAE pre-training strategy to a hybrid state-of-the-art (SOTA) medical image segmentation architecture, UNETR++ [18], and co-pretrain UNETR++ alongside the ViT. Specifically, masked crops are processed separately by both ViT and UNETR++ to reconstruct the associated missing patches. Reconstruction consistency loss and spatial consistency loss (derived from the pre-text task) are employed to connect the two different types of architectures in pre-training, enhancing their representation learning capability.
@@
Our main contributions are summarized as follows: (1) We propose a new topological loss and introduce a pre-text task for MAEs to learn geometric shape and spatial information. (2) We extend the MAE pre-training strategy to a hybrid SOTA medical image segmentation architecture and co-pretrain it alongside ViT. (3) We develop a fine-tuned model for downstream segmentation tasks, and demonstrate the effectiveness of our new approach on five public 3D segmentation datasets.
@@
Figure 1: Illustrating the effect of our proposed topological loss. Left: raw image examples of the Synapse CT dataset; middle: reconstructed images with the mean squared error (MSE) loss [28]; right: reconstructed images with a combination of the MSE and proposed topological losses.
Figure 2: An overview of our proposed pipeline.
II Method
@@
II-A Capturing the Topology of Input Volumes
Given a 3D image $I$, we represent $I$ with a cubical complex $C$. Typically, the cubical complex $C$ takes each voxel of $I$ as an individual vertex and contains connectivity information on vertex neighbourhoods via edges, squares, and their higher-dimensional counterparts [14, 8]. In this work, we use persistent homology (PH) [7] to extract topological features of different dimensions from $C$, including connected components (0-D), cycles/loops ($1$-D), and voids ($2$-D). PH combines the homology of super-level sets by sweeping a threshold function through the entire real numbers. Specifically, for a threshold value $\tau\in\mathbb{R}$, a cubical complex is defined as: $C^{(\tau)}:=\{x\in I\ |\ f(x)\geq\tau\},$ where $f(x)$ is the voxel value of $x$. When sweeping the threshold, the topology changes only at a finite number of values, $\tau_{1}\geq\tau_{2}\geq\cdots\geq\tau_{m-1}\geq\tau_{m}$, and we obtain a sequence of nested cubical complexes, $\emptyset\subseteq C^{(\tau_{1})}\subseteq C^{(\tau_{2})}\subseteq\cdots\subseteq C^{(\tau_{m-1})}\subseteq C^{(\tau_{m})}=I$, which forms the super-level set filtration. PH tracks topological features across all the complexes in this filtration, representing each feature as a tuple $(\tau_{i},\tau_{j})$ with $\tau_{i}\geq\tau_{j},$ indicating the cubical complex in which a feature appears and disappears, respectively. For example, a 0-D tuple $(\tau_{i},\tau_{j})$ represents a connected component that appears at threshold $\tau_{i}$ and disappears at threshold $\tau_{j}$. The tuples of the $k$-D ($0\leq k\leq 2$) features are saved in the $k$-th persistence diagram $D_{I}^{k}$, which is a multi-scale shape descriptor of all topological features of the 3D image $I$.
[…]
Constructing Topological Loss. Given an input volume $I$ and a reconstructed volume $I^{\prime}$, our new topology-aware loss is defined as: $\mathcal{L}_{topo}(I,I^{\prime})=\left(\sum_{i=0}^{2}(W_{2}(D_{I}^{i},D_{I^{\prime}}^{i}))^{2}\right)^{\frac{1}{2}},$ where $W_{2}(\cdot,\cdot)$ denotes the 2-Wasserstein distance, and $D_{I}^{i}$ and $D_{I^{\prime}}^{i}$ are the $i$-D persistence diagrams of $I$ and $I^{\prime}$, respectively. Note that our proposed topological loss differs from that in [19]. First, we extract topological features from 3D images, not from the segmentation. Second, we use the topological loss for self pre-training and medical image segmentation, not for 3D reconstruction tasks.
II-B Exploiting Global Spatial Information
@@
Constructing Spatial Loss. We denote the 9 key points of a 3D crop as $(p_{1},p_{2}$, $\ldots,p_{9})$, where $p_{9}$ is for the crop center, and each $p_{i}=(x_{i},y_{i},z_{i})$. Given the ground truth (GT) and the prediction of the 9 positions, $P=(p_{1},p_{2},\ldots,p_{9})$ and $P^{\prime}=(p^{\prime}_{1},p^{\prime}_{2},\ldots,p^{\prime}_{9})$, the spatial loss is defined as: $\mathcal{L}_{spa}(P,P^{\prime})=\mathcal{L}_{MSE}(P,P^{\prime}),$ where $\mathcal{L}_{MSE}$ is the MSE loss. Our spatial loss definition and implementation are different from those in [23].
II-C Co-pretraining the ViT and UNETR++ Models
@@
Constructing Spatial Consistency Loss. Given predictions of 9 key point positions from the ViT and UNETR++ encoders, $P^{\prime}=(p^{\prime}_{1},p^{\prime}_{2},\ldots,p^{\prime}_{9})$ and $P^{\prime\prime}=(p^{\prime\prime}_{1},p^{\prime\prime}_{2},\ldots,p^{\prime\prime}_{9})$, the spatial consistency loss is defined as: $\mathcal{L}_{spa-consis}(P^{\prime},P^{\prime\prime})=\mathcal{L}_{MSE}(P^{\prime},P^{\prime\prime}).$ The spatial consistency loss aligns the reconstructed images at the spatial level, enhancing the feature learning capability of both the ViT and UNETR++ models.
Constructing Reconstruction Consistency Loss. Given reconstructed volumes $I^{\prime}$ and $I^{\prime\prime}$ by the ViT and UNETR++ models, our reconstruction consistency loss function computes the MSE between the reconstructed volumes $I^{\prime}$ and $I^{\prime\prime}$, as: $\mathcal{L}_{rec-consis}(I^{\prime},I^{\prime\prime})=\mathcal{L}_{MSE}(I^{\prime},I^{\prime\prime}).$ We compute this loss only on masked patches, similar to MAE in [11]. The reconstruction consistency loss aligns the reconstructed images at the image level, further enhancing the representation learning capability of both ViT and UNETR++.
@@
II-D Constructing the Fine-tuned Architecture
@@
TABLE I: Segmentation results on the Synapse CT dataset. The best results are marked in bold, and the second-best results are underlined.
@@
MAE + UNETR++ (ours) 85.96M 82.49 95.68 89.30 87.64 79.60 96.98 88.47 92.58 81.27 88.94 $(1.72\%\uparrow)$ 5.89 $(1.64\uparrow)$
$p$-values <5e2<5e-2 (Dice), <1e2<1e-2 (HD95)
III Experiments and analysis
III-A Datasets and Experimental Setup